# VisualCoder: Guiding Large Language Models in Code Execution with Fine-grained Multimodal Chain-of-Thought Reasoning

### v2

@@ 5.2 CFG Images vs Text-Based Descriptions
| Claude-3.5-Sonet | 60.5 | 74.0 |
| Gemini-1.5-Flash | 65.3 | 74.1 |
@@
To assess the impact of visual representations in coding tasks, specifically in Code Execution Prediction, we conducted an experiment where LLM models were provided with either Mermaid-format (text-based) (see one example of Mermaid-format in A.1) or image-based CFGs, along with the input, and tasked with predicting the code’s output. The prompt remained the same as the previous experiment, but the models received CFGs instead of code. Results in Table 2 show that CFG images significantly boost performance in reasoning tasks, underscoring the value of visual aids in enhancing Multimodal LLMs’ reasoning.
@@ 5.3 VisualCoder Multi-modal Reasoning
Experimental Setting. This experiment involved two tasks: Program Repair and Fault Localization (further details are provided in A.3). For Program Repair, we generated a dataset from LiveCodeBench Jain et al. (2024), selecting 400 instances to avoid the saturation seen in simpler benchmarks like MBPP-S Odena et al. (2021). We sampled six solutions using Claude 3.5 Sonnet and Haiku with a 3:1 ratio, ensuring varied difficulty levels. After filtering out fully correct or incorrect solutions, we ensure the direction of code is correct but with some mistakes. We finalized 384 solutions for 173 problems. For Fault Localization, we used the FixEval dataset Haque et al. (2022), which has about 210 programs with diverse runtime errors. The prompt used was listed in A.4.
@@
Experimental Results. Table 3 provides a detailed comparison of VisualCoder with other baseline methods across multiple settings. Chain-of-Thought (CoT) reasoning generally improved model performance, as seen in the Program Repair task where Claude Sonnet 3.5 improved from 64.1% (plain code) to 63.0% (with CoT), and GPT-4o increased from 38.7% to 40.1%. However, combining CoT with CFG images caused a notable performance drop across all models. For instance, Claude Sonnet 3.5’s accuracy dropped from 63.0% to 55.5%, GPT-4o fell from 40.1% to 37.6%, and InternVL2-26B dropped from 4.0% to 2.1%. Similar declines occurred in the Fault Localization task. This suggests that while CFGs offer structural insights, integrating them with CoT without proper schemes can confuse the models and reduce accuracy, a finding consistent with Zhang et al. (2023).
In the Program Repair task, which relies more on logical reasoning than execution-heavy tasks, CFGs proved less useful. Although our method didn’t outperform the highest-performing settings (e.g., plain code without CoT for Claude at 64.1% or GPT-4o at 40.1%), it significantly boosted performance over the plain code + CFG w/ CoT setting. It raised Claude Sonnet 3.5 from 55.5% to 62.9%, and GPT-4o from 37.6% to 38.7%. InternVL2-26B, which struggled with CFG + CoT (2.1%), improved to 6.3% with VisualCoder and 10.7% when combined with Multimodal-CoT. In some cases, it outperformed methods like NeXT and Multimodal-CoT, with Claude Sonnet 3.5 achieving 62.9% with VisualCoder, compared to 57.3% with NeXT and 58.7% with Multimodal-CoT, showing its capability, even in tasks where CFGs are less central.
@@
| Plain code w/ CoT | 63.0 | 38.9 | 4.0 | |
| Plain code + CFG w/o CoT | 61.2 | 36.7 | 0.9 | |
| Program Repair | Plain code + CFG w/ CoT | 55.5 | 40.1 | 2.1 |
@@
The real improvement came with VisualCoder. As shown in Table 3, VisualCoder achieved the highest accuracy for both Claude Sonnet 3.5 (91.4%) and GPT-4o (90.4%). When combined with Multimodal-CoT, performance further improved, with Claude reaching 92.8% and GPT-4o 91.9%. The biggest gain was for InternVL2-26B, which increased from 41.0% (CoT with CFG) to 53.6% with VisualCoder and Multimodal-CoT. These results show that integrating CFGs with CoT reasoning and the Reference Mechanism significantly boosts fault localization, especially when paired with Multimodal-CoT. This also highlights the effectiveness of generating rationale that efficiently leverages both plain code and CFG images.
@@ 6 Qualitative Analysis
The first three rows of Figure 2 show Claude Sonnet 3.5’s outputs under different prompt settings, all failing to fully grasp the code’s complexity. In the left example (a use-before-initialization error), the model wrongly identifies lst[0] as the issue, missing the control flow dependencies affecting lst’s initialization. In the right example (unreachable code), it highlights G += 1 but overlooks the actual problem: using a float N in the range function. These errors highlight the limitations of plain code reasoning, even with CFG or CoT.
@@ 7 Attention Pattern Analysis
As shown in Figure 3, the attention maps reveal key differences between the Reference CoT and vanilla CoT approaches. With the Reference CoT, the InterVL2-26B model consistently attends to the relevant nodes in the CFG image at each reasoning step. In contrast, the vanilla approach exhibits a more diffuse attention pattern, occasionally focusing on irrelevant regions, which could contribute to poorer performance in debugging tasks. Figure 4 also shows that with Reference Mechanism, VLM more focuses on vision tokens in CFG image for reasoning, leading to better capturing of the alignment between code and CFG image.
These findings also corroborate our intuition behind Equation (5). By incorporating reference mapping, the model adopts a “more focused” attention mechanism for each Chain-of-Thought (CoT) step, facilitating more precise reasoning.
@@ 8 Conclusion
In conclusion, VisualCoder enhances LLMs’ understanding of code execution by integrating multimodal inputs, specifically control flow graph (CFG) visualizations. While traditional LLMs handle static code well, they struggle with dynamic behaviors. Our Reference CoT technique links code lines with CFG elements, reducing errors and providing a clearer representation of execution flow. Our results show that visual CFGs improve LLMs’ reasoning capabilities on code execution.
@@ 9 Limitations
Our method has not yet been extensively tested on large-scale codebases or repository-level projects with numerous files. While it performs well on individual code snippets, generating a full CFG for an entire repository presents significant challenges due to the complexity and scale involved. Constructing CFGs for large repositories with many interconnected files can be computationally expensive and difficult to manage efficiently.
@@
Finally, our method has been primarily evaluated on Python-based benchmarks, leaving its generalizability across other programming languages in question. Since different languages have distinct control structures and syntax, VisualCoder’s ability to generate accurate CFGs and reason effectively may vary. Future work should explore its applicability to a wider range of programming languages and paradigms to assess its broader utility.

### v3

@@ 5.2 CFG Images vs Text-Based Descriptions
| Claude 3.5 Sonet | 60.5 | 74.0 |
| Gemini 1.5 Flash | 65.3 | 74.1 |
@@
To assess the impact of visual representations in coding tasks, specifically in Code Execution Prediction, we conducted an experiment where LLM models were provided with either Mermaid-format (text-based) (see one example of Mermaid-format in A.1) or image-based CFGs, along with the input, and tasked with predicting the code’s output. The prompt remained the same as the previous experiment, but the models received CFGs as images instead of as texts. Results in Table 2 show that CFG images significantly boost performance in reasoning tasks, underscoring the value of visual aids in enhancing Multimodal LLMs’ reasoning.
@@ 5.3 VisualCoder Multi-modal Reasoning
Experimental Setting. This experiment involved two tasks: Program Repair and Fault Localization (further details are provided in Section A.3). For Program Repair, we generated a dataset from LiveCodeBench Jain et al. (2024), selecting 400 instances to avoid the saturation seen in simpler benchmarks like MBPP-S Austin et al. (2021). We sampled six solutions for each instance using Claude 3.5 Sonnet and Haiku with a 3:1 ratio, ensuring varied difficulty levels. After filtering out fully correct solutions (those passing all test cases) and completely incorrect ones (those failing all test cases), we retained solutions that have the correct direction—passing a subset of test cases but containing some errors. We finalized 384 solutions for 173 problems. For Fault Localization, we used the FixEval dataset Anjum Haque et al. (2023), which has about 210 programs with various runtime errors. The prompt used was listed in Section A.4.
Unlike the previous sections, we selected models with stronger code reasoning abilities to handle the increased complexity of Program Repair and Fault Localization task. Specifically, Claude 3.5 Sonnet was retained for its robust performance, GPT-4o replaced Gemini 1.5 Flash due to its superior capabilities and wider adoption, and InterVL2-26B replaced its 8B version, which struggled with coding tasks, often producing generic or incorrect answers.
@@
Experimental Results. Table 3 provides a detailed comparison of VisualCoder with other baseline methods across multiple settings. Chain-of-Thought (CoT) reasoning generally improved model performance, as seen in the Program Repair task where GPT-4o improved from 38.7% (plain code) to 40.1% (with CoT), and InternVL2 increased from 0.4% to 4.0%. However, combining CoT with CFG images caused a notable performance drop across all models. For instance, Claude 3.5 Sonnet’s accuracy dropped from 63.0% to 55.5%, GPT-4o fell from 40.1% to 37.6%, and InternVL2-26B dropped from 4.0% to 2.1%. Similar declines occurred in the Fault Localization task. This suggests that while CFGs offer structural insights, integrating them with CoT without proper schemes can confuse the models and reduce accuracy, a finding consistent with Zhang et al. (2023).
In the Program Repair task, which relies more on logical reasoning than execution-heavy tasks, CFGs proved less useful. Although our method didn’t outperform the highest-performing settings (e.g., plain code without CoT for Claude at 64.1%), it significantly boosted performance over the plain code + CFG w/ CoT setting. It raised Claude 3.5 Sonnet from 55.5% to 62.9%, and GPT-4o from 37.6% to 41.2%. InternVL2-26B, which struggled with CFG + CoT (2.1%), improved to 6.3% with VisualCoder and 10.7% when combined with Multimodal-CoT. In some cases, it outperformed methods like NeXT and Multimodal-CoT, with Claude 3.5 Sonnet achieving 62.9% with VisualCoder, compared to 57.3% with NeXT and 58.7% with Multimodal-CoT, showing its capability, even in tasks where CFGs are less central.
@@
| Plain code w/ CoT | 63.0 | 40.1 | 4.0 | |
| Plain code + CFG w/o CoT | 61.2 | 36.5 | 0.9 | |
| Program Repair | Plain code + CFG w/ CoT | 55.5 | 37.6 | 2.1 |
@@
As shown in Table 3, VisualCoder achieved the highest accuracy for both Claude 3.5 Sonnet (91.4%) and GPT-4o (90.4%). When combined with Multimodal-CoT, performance further improved, with Claude reaching 92.8% and GPT-4o 91.9%. The biggest gain was for InternVL2-26B, which increased from 41.0% (CoT with CFG) to 53.6% with VisualCoder and Multimodal-CoT. These results show that integrating CFGs with CoT reasoning and the Reference Mechanism boosts fault localization, especially when paired with Multimodal-CoT. This also highlights the effectiveness of generating rationale that efficiently leverages both plain code and CFG images.
VisualCoder with Complex Java Repositories. To further evaluate VisualCoder ’s effectiveness on more complex, real-world software defects, we extended our experiments to the Defects4J v1.0 Just et al. (2014) benchmark. Defects4J comprises 245 real-world Java bugs from five open-source projects: Chart, Closure, Lang, Math, and Time. For this setting, we used GPT-4o and evaluated its fault localization performance using the acc@k metric, which quantifies the number of bugs where the actual buggy location appears among the top k predictions generated by a tool. Table 4 presents the results, comparing Plain Code (Vanilla), Plain Code + CFG, and Plain Code + CFG + Reference Mechanism as VisualCoder.
| Model | acc@1 | acc@4 | acc@10 |
| Plain code | 47 | 73 | 90 |
| Plain code + CFG | 54 | 78 | 95 |
| VisualCoder | 59 | 80 | 97 |
As shown in Table 4, VisualCoder consistently outperforms the baselines across all metrics, demonstrating its effectiveness in complex fault localization tasks. The integration of CFGs and the Reference Mechanism significantly enhances localization accuracy, particularly at acc@1, where VisualCoder correctly identifies 12 more bugs than the baseline. These results highlight VisualCoder ’s robustness in large-scale Java projects, reinforcing its potential for software debugging.
@@ 6 Qualitative Analysis
The first three rows of Figure 2 show Claude Sonnet 3.5’s outputs under different prompt settings, all failing to fully grasp the code’s complexity. In the left example (a use-before-initialization error), the model incorrectly identifies lst[0] as the issue, missing the control flow dependencies affecting lst’s initialization. In the right example (unreachable code), it highlights G += 1 but overlooks the actual problem: using a float N in the range function. These errors highlight the limitations of plain code reasoning, even with CFG or CoT.
@@ 7 Attention Pattern Analysis
As shown in Figure 3, the attention maps reveal key differences between the Reference CoT and vanilla CoT approaches. With the Reference CoT, the InterVL2-26B model consistently attends to the relevant nodes in the CFG image at each reasoning step. In contrast, the vanilla approach exhibits a more diffuse attention pattern, occasionally focusing on irrelevant regions, which could contribute to poorer performance in debugging tasks. Figure 4 also shows that with Reference Mechanism, VLM more focuses on vision tokens in CFG image for reasoning, leading to better capturing of the alignment between code and CFG image. These findings also corroborate our intuition behind Equation 5. By incorporating reference mapping, the model adopts a “more focused” attention mechanism for each Chain-of-Thought step, facilitating more precise reasoning on program execution.
@@ 8 Conclusion
In conclusion, VisualCoder enhances LLMs’ reasoning about code execution by incorporating multimodal inputs, specifically control flow graph (CFG) visualizations. Traditional LLMs, while effective at processing static code syntax, struggle to capture dynamic execution behaviors, leading to incorrect predictions and limited reasoning about program flow. By introducing the Reference CoT technique, VisualCoder establishes explicit connections between source code lines and CFG elements, ensuring a structured and interpretable representation of execution logic. This approach reduces reasoning errors, improves alignment between textual and visual execution cues, and enables more accurate program behavior predictions. Our experimental results show that augmenting LLMs with visual CFGs significantly improves performance over text-based CFG descriptions alone, validating our multimodal approach.
Acknowledgments
Tien N. Nguyen was supported in part by the US National Science Foundation (NSF) grant CNS-2120386 and the National Security Agency (NSA) grant NCAE-C-002-2021.